OpenStudy (anonymous):

Solve. 7^(x – 5) = 7^(4x + 7)

OpenStudy (anonymous):

\[ \huge 7x-5=4x+7 \] \[\huge 7x-4x = 7+5 =12\] \[\huge 3x =12\] \[\huge x=\frac { 12 }{ 3 } =4 \]

OpenStudy (anonymous):

Thank you for showing how to do it!!

OpenStudy (anonymous):

lol my pleasure

sam (.sam.):

The question should be \[\huge 7^{x-5}=7^{4 x+7}\] \[\huge \begin{array}{l} \mathbb{Same~ base} \\ x-5=4 x+7 \\ \text{Add }(5-4 x)\text{ both sides} \\ -3 x=12 \\ \text{Divide both sides by }-3 \\ x=-4 \\\end{array}\]